OpenStudy (anonymous):

Which math operation determines the following pattern? 1, 4, 16, 64, ...? addition subtraction multiplication division Which math operation determines the following pattern? 3, 7, 10, 17, ... addition subtraction multiplication division

OpenStudy (anonymous):

@amistre64 @ajprincess help?

OpenStudy (anonymous):

@KeganMoore @eSpeX

OpenStudy (anonymous):

64/4=16 16/4=4 4/4=1 So the first one is division

OpenStudy (anonymous):

and the 2nd?

OpenStudy (anonymous):

I am thinking, the second one is tricky... hold on.

OpenStudy (espex):

Look at the pattern, in one you have a term that plays off the product of the two before it and the other is the sum.

OpenStudy (anonymous):

I can't figure out the second one. :/ sorry

OpenStudy (espex):

In the second one, what relationship can you find with the first three terms?

OpenStudy (anonymous):

For the second one, you add the preceding two terms to get the next term. 3+7=10 7+10=17....

OpenStudy (anonymous):

So it's addition.